In a recent year, Delaware had the highest per capita annual income with $51,803. If standard deviation= $4850, what is the probability that a random sample of 36 state residents had a mean income greater than $50,100
Answer #1

Solution :

Given that ,

= / n = 4850 / 36 = 808.3333

P( > 50100) = 1 - P( < 50100)

= 1 - P[( - ) / < (50100 - 51803) / 808.3333]

= 1 - P(z < -2.11)

= 1 - 0.0174

= 0.9826

Probability = 0.9826
Answer #1

Solution :

Given that ,

= / n = 4850 / 36 = 808.3333

P( > 50100) = 1 - P( < 50100)

= 1 - P[( - ) / < (50100 - 51803) / 808.3333]

= 1 - P(z < -2.11)

= 1 - 0.0174

= 0.9826

Probability = 0.9826